# Compiler Optimization via LLM Reasoning for Efficient Model Serving

### v1

@@ 4 Results
We implement the Reasoning Compiler as an extension to MetaSchedule [5]. The framework introduces three modular components: (1) a prompt generator that serializes the current scheduling state, including the IRModule, transformation trace (i.e., the applied schedule history), and hardware cost model outputs, into structured prompts that capture the textual difference from the base IRModule and reflect the current schedule’s performance; (2) an LLM interface that queries an external API (e.g., OpenAI) and parses LLM output into candidate transformation sequences; and (3) a tree manager that performs MCTS with selection based on UCT score, expansion using LLM suggested transformations, simulation with hardware-aware cost model, and backpropagation for tree statistics updates.
@@ 4.1 Experimental Setup
We evaluate Reasoning Compiler on four representative computational kernels from production-scale models: (1) Self-Attention layer from Llama3-8B [10], (2) Mixture-of-Experts (MoE) layer from DeepSeek-R1 [11], (3) Self-Attention layer from FLUX (Stable Diffusion) [12], and (4) Convolution layer from FLUX [12]. Compiler optimization is framed as a sequential decision process and guided by MCTS [13], using the Upper Confidence Bound for Trees (UCT) criterion [8] with exploration parameter and a branching factor , following prior work [14, 15]. During search, the LLM (OpenAI GPT-4o mini [16]) is queried using hierarchical context—specifically, the parent and grandparent schedules and their transformations—to enable informed proposal generation. We compare three optimization strategies: (1) MetaSchedule [5], which uses Evolutionary Search; (2) MCTS without LLM involvement (MCTS); and (3) LLM+MCTS that uses prompt-based proposal generation (LLM-Guided MCTS). All experiments are conducted on a dedicated Intel Core i9 CPU using Apache TVM v0.20.0 [6] [17], with the hardware environment held constant to isolate the effects of scheduling decisions. Each experiment is repeated 20 times, and we report the mean performance to ensure statistical stability. We further eliminate noise by disabling background processes and ensuring no competing workloads during measurement. Additionally, we leverage OpenAI and HuggingFace model serving APIs to access the respective models. The implementation is open-sourced.
@@ 4.2 Evaluation
We assess the sample efficiency of our LLM-guided compilation framework by analyzing how code quality evolves with increasing search budget, quantified in terms of evaluated transformation proposals. Figure 3 and Figure 4 present results across four representative workloads, encompassing both transformer-style attention layers and convolution-heavy architectures. Across all benchmarks, our method achieves competitive or superior code performance with significantly fewer samples than state-of-the-art black-box autotuners such as MetaSchedule with Evolutionary Search. These results directly support the central hypothesis of our work: that leveraging LLM-driven, context-aware reasoning enables more efficient and effective exploration of the compiler optimization space.
@@
Rapid convergence in low-sample regimes. A consistent trend across all benchmarks is the rapid ascent of code quality in the initial stages of search. This early-stage performance is critical in practice, as real-world compiler pipelines often operate under strict tuning time budgets. Figure 3 shows Relative Speedup over Pre-Optimized Code on the y-axis, with the number of evaluated transformation proposals on the x-axis. Speedup is defined as the ratio of the execution time of the unoptimized code to that of the optimized code after tuning. Higher values indicate more efficient and optimized codes. For instance, on the Llama3-8B Self-Attention Layer, the proposed method achieves a 7.08 speedup over the untuned baseline with just 36 samples, whereas Evolutionary Search requires 72 samples, which is twice the budget to achieve comparable gains. On the DeepSeek-R1 MoE Layer, the gap is even more pronounced: LLM-guided search achieves 7.05 speedup at 36 samples, while Evolutionary Search falls short of this mark even after 3000 samples.
@@
Quantitative sample efficiency. To formally quantify sample efficiency, we compare the number of samples required by each method to reach target speedups. On the FLUX Self-Attention Layer, LLM-Guided MCTS attains a 2 speedup using only 36 samples, while Evolutionary Search requires more than 600 samples—a 16 reduction in tuning cost. On the FLUX Convolution Layer, our approach consistently outperforms Evolutionary Search across nearly all budget levels and reaches Evolutionary Search ’s final performance after evaluating just 400 samples.
@@
Speedup relative to baselines. Figure 4 shows Relative Speedup over Evolutionary Search-Generated Code, again with number of samples on the x-axis. Here, speedup is the ratio between Evolutionary Search ’s best code and ours at each point in the search. Values greater than 1 indicate superior performance relative to Evolutionary Search. These metrics highlight that LLM-Guided MCTS not only produces better codes, but does so more aggressively and earlier in the search process. For example, on the DeepSeek-R1 MoE Layer, LLM-Guided MCTS achieves a 3.3 speedup over Evolutionary Search at 36 samples, with the gap narrowing over time as Evolutionary Search eventually converges. This trend, which shows strong initial gains followed by convergence, demonstrates that LLM-Guided MCTS quickly identifies high-performing regions of the search space, while Evolutionary Search ’s uninformed search requires substantial exploration to reach similar quality.
@@
Operator-specific trends. We observe that certain operator types, such as matrix multiplication operations extracted from attention layers, exhibit sharper performance improvements. This is likely due to recurring structural patterns such as loop fusion, tiling, and vectorization, which pretrained LLMs can more readily recognize and exploit. Convolutional operators, by contrast, expose a broader and less regular transformation space. Nonetheless, our approach consistently matches or exceeds baseline performance with fewer samples, underscoring its effectiveness across diverse operator characteristics.
@@
Implications. These findings reinforce our core thesis: compiler optimization should be cast as a structured decision process, enriched by prior knowledge and contextual reasoning. Our integration of LLMs into Monte Carlo tree search results in a strategically guided and sample-efficient search, particularly valuable in scenarios with constrained tuning budgets. By generating performant codes with orders-of-magnitude fewer samples, our framework offers both practical deployment advantages and a compelling alternative to conventional, sample-inefficient compilation pipelines.
@@ 4.3.1 Impact of LLM Choice and Reasoning Strategy
To better understand the contributions of different components in our approach, we conduct an ablation study focused on the effects of LLM selection and reasoning modality. Figure 5(a) shows the relative speedup over unoptimized code as a function of the number of schedule samples evaluated by LLM-Guided MCTS on the Llama3-8B Self-Attention Layer benchmark, using a range of LLM models for API calls. The x-axis indicates the cumulative number of schedules explored, while the y-axis shows the best speedup achieved so far. This setup enables us to directly compare how effectively various LLMs leverage contextual information to guide the search. These general trends of the results support our central claim: compiler optimization benefits from goal-directed, context-aware reasoning in terms of sample efficiency. Below, we discuss the specific behaviors that exemplify different reasoning strategies.
Large instruction-tuned Llama3.3 (70B) achieves exceptional sample efficiency. The instruction-tuned Llama3.3-70B model rapidly attains near-optimal performance, reaching a 9.69 speedup after only 36 samples—roughly 86% of the GPT-4o mini’s maximum speedup but using less than 6% of its sampling budget. This corresponds to an approximately 15 improvement in sample efficiency. Instruction tuning also significantly improves the ability of LLMs to generate domain-specific, context-aware transformation proposals. The consistent performance advantage of instruction-tuned models over untuned counterparts of comparable size confirms that semantic task alignment, combined with sufficient model capacity, synergistically enhances the effectiveness of sequential context reasoning in guiding compiler optimizations.
@@
Lower-parameter models also achieve high sample efficiency. Despite their reduced scale, smaller models still produce notable speedups relative to the untuned baseline. For example, Llama3.1-Instruct (8B) reaches a 5.87 speedup, and DeepSeek-R1-Distill-Qwen (7B) achieves 4.86 at just 36 samples. When compared to the widely used Evolutionary Search strategy, which requires around 72 samples to achieve a 7 speedup and fails to reach comparable performance for tuning the DeepSeek-R1 MoE Layer even after 3000 samples, these smaller models consistently outperform. LLM-Guided MCTS with lower-parameter models achieve at least twice the sample efficiency of Evolutionary Search, making them well-suited for efficient compiler optimization in local or edge deployments.
Open-source models match proprietary model performance. Our results demonstrate that open-source LLMs, when adequately scaled and instruction-tuned, match or exceed the performance of proprietary baselines such as GPT-4o mini. This underscores the broad applicability of our approach and its independence from proprietary data or architectures, enabling widespread adoption of context-aware, LLM-guided compiler optimization.
@@ 4.3.2 Impact of Historical Trace Depth on Optimization Efficiency
Figure 5(b) presents the relative speedup over unoptimized code as a function of the number of schedule samples evaluated by LLM-Guided MCTS on the Llama3-8B Self-Attention Layer benchmark. Using a deeper historical trace (see Figure 1) in the prompt (parent + grandparent + great-grandparent) leads to faster convergence compared to the shallower trace (parent + grandparent). For example, at 36 samples, the deeper trace achieves a speedup of approximately 7.13, slightly surpassing the 7.08 of the shallower trace. However, by 72 samples, the deeper trace saturates at 11.36 speedup, while the shallower trace reaches only 8.38, requiring many more samples (around 579) to approach 11.3 performance. This demonstrates that including longer historical context enables the LLM to better capture dependencies and synergies in transformation sequences, resulting in more sample-efficient and goal-directed exploration, validating the advantage of context-aware reasoning.
@@ 5 Related Work
ML-Based Auto-Tuning. Auto-tuning frameworks optimize performance-critical parameters (e.g., loop tile sizes, phase orderings, memory layouts) using a variety of ML-based techniques, including linear models [18, 19], tree-based methods [20, 21], Bayesian networks [22, 23], evolutionary algorithms [20, 24, 25], clustering [19, 25], and reinforcement learning [26, 24, 7, 25]. The Reasoning Compiler shares the same goal of performance-driven parameter selection, but distinguishes itself by combining LLM-based contextual reasoning with structured search (via MCTS) to explore transformation sequences in a history- and structure-aware manner.
Machine Learning for Neural Compilation. Machine learning has been extensively applied to optimize neural network inference pipelines, including high-level graph-level optimizations [27, 28, 29, 30, 31, 32, 33, 34] and low-level code generation [3, 35, 36, 37, 6, 38, 39, 40]. Systems such as TVM/Ansor [6, 2, 4] and FlexTensor [41] employ learned cost models and evolutionary strategies to navigate large configuration spaces. While these approaches are highly effective at tuning tensor programs, they typically focus on local parameter optimization or rely on domain-specific heuristics. The Reasoning Compiler moves beyond these works by introducing contextual reasoning through LLMs, enabling the system to reason over transformation history, structural changes, and performance trends, an approach not explored in prior neural compilation work.
LLMs for Code Reasoning and Optimization. LLMs have demonstrated capabilities in code generation [42, 43, 44, 45, 46, 47], fuzzing [48], and bug repair [49]. Recent work has explored the use of LLMs to generate phase orderings or perform disassembly [50, 51]. The Reasoning Compiler advances these approaches by embedding an LLM in a structured decision loop, leveraging it for context-aware reasoning within a grounded search process.

### v3

@@ 4 Results
We implement Reasoning Compiler as an extension to MetaSchedule [metaschedule:neurips:2022]. The framework introduces three modular components: (1) a prompt generator that serializes the current scheduling state, including the IRModule, transformation trace (i.e., the applied schedule history), and hardware cost model outputs, into structured prompts that capture the textual difference from the base IRModule and reflect the current schedule’s performance; (2) an LLM interface that queries an external API (e.g., OpenAI) and parses the LLM’s output into candidate transformation sequences; and (3) a tree manager that performs MCTS with selection based on UCT score, expansion using LLM suggested transformations, simulation with a hardware-informed cost model, and backpropagation for tree statistics updates.
@@ 4.1 Experimental Setup
We evaluate Reasoning Compiler on five representative computational kernels drawn from production-scale models: (1) self-attention layer from Llama-3-8B [llama:arxiv:2024], (2) mixture-of-experts (MoE) layer from DeepSeek-R1 [deepseek:arxiv:2025], (3) self-attention layer from FLUX (stable diffusion) [flux:gitrepo:2024], (4) convolution layer from FLUX [flux:gitrepo:2024], and (5) MLP Layer from Llama-4-Scout [llama4:huggingface:2025]. Compiler optimization is framed as a sequential decision process and guided by MCTS [mcts], using the Upper Confidence bounds applied to Trees (UCT) criterion [uct:ecml:2006] with exploration parameter and branching factor , following prior work [branching1:cg:2007, ucb:ml:2002]. During search, the LLM (OpenAI GPT-4o mini [openai:4o-mini:2025]) is queried using hierarchical context—specifically, the parent and grandparent schedules and their transformations—to enable informed proposal generation. We compare three optimization strategies: (1) TVM MetaSchedule [metaschedule:neurips:2022], which uses Evolutionary Search; (2) MCTS without LLM guidance (MCTS); and (3) Reasoning Compiler that uses prompt-based proposal generation (LLM-Guided MCTS). All experiments are conducted using Apache TVM v0.20.0 [tvm, tvm:v0200:2025]. Our main experimental environment is a dedicated Intel Core i9 workstation under a fixed software and hardware stack to isolate scheduling effects. This main environment covers all five kernels above, and is the ablation environment. To show portability and scalability across consumer and datacenter processors, we evaluate each of the five kernels on a total of five hardware platforms: Amazon Graviton2, AMD EPYC 7R13, Apple M2 Pro, Intel Core i9, and Intel Xeon E3. We also report end‑to‑end Llama-3-8B results across the same five platforms. Each experiment is repeated 20 times, and we report the mean performance to ensure statistical stability. We further eliminate noise by disabling background processes and ensuring no competing workloads during measurement. Additionally, we leverage OpenAI and HuggingFace model serving APIs to access the respective models. The implementation is open-sourced.
@@ 4.2 Evaluation
We assess the sample efficiency of our LLM-guided compilation framework by analyzing how code quality evolves with increasing search budget, quantified in terms of evaluated transformation proposals. Figure 3 presents results across five representative workloads, encompassing both transformer-style attention layers and convolution-heavy architectures. Across all benchmarks, our method achieves competitive or superior code performance with significantly fewer samples than state-of-the-art black-box autotuners such as MetaSchedule with Evolutionary Search. These results directly support the central hypothesis of our work: leveraging LLM-driven, context-aware reasoning enables more efficient and effective exploration of the compiler optimization space.
@@
Rapid convergence in low-sample regimes. A consistent trend across all benchmarks is the rapid ascent of code quality in the initial stages of search. This early-stage performance is critical in practice, as real-world compiler pipelines often operate under strict tuning time budgets. Figure 3 shows Relative Speedup over Pre-Optimized Code on the y-axis, with the number of evaluated transformation proposals on the x-axis. Speedup is defined as the ratio of the execution time of the unoptimized code to that of the optimized code after tuning. Higher values indicate more efficient and optimized code. For instance, on the Llama-3-8B Attention Layer, LLM-Guided MCTS achieves a 7.08 speedup over the untuned baseline with just 36 samples, whereas Evolutionary Search requires 72 samples, which is twice the budget to achieve comparable gains. On the Llama-4-Scout MLP Layer, the gap is even more pronounced: LLM-Guided MCTS achieves 12.7 speedup at 20 samples, while Evolutionary Search falls short of this mark even after 3000 samples.
@@
Quantitative sample efficiency. To formally quantify sample efficiency, we compare the number of samples required by each method to reach target speedups. On the FLUX Attention Layer, LLM-Guided MCTS attains a 2 speedup using only 36 samples, while Evolutionary Search requires more than 600 samples, a 16 reduction in tuning cost. On the FLUX Convolution Layer, LLM-Guided MCTS consistently outperforms Evolutionary Search across nearly all budget levels and reaches Evolutionary Search ’s final performance after evaluating just 400 samples.
| Hardware Platform | Benchmark | TVM | Reasoning Compiler | Improvement | |||
| # Samples | Speedup | # Samples | Speedup | Sample Reduction | Sample Efficiency Gain | ||
| Amazon Graviton2 | Llama-3-8B Attention Layer | 510 | 3.9 | 60 | 5.1 | 8.5 | 11.1 |
| Deepseek-R1 MoE Layer | 980 | 2.7 | 150 | 5.9 | 6.5 | 14.4 | |
| FLUX Attention Layer | 320 | 1.6 | 130 | 3.0 | 2.5 | 4.6 | |
| FLUX Convolution Layer | 160 | 1.8 | 20 | 4.1 | 8.0 | 18.2 | |
| Llama-4-Scout MLP Layer | 1,630 | 1.7 | 500 | 4.0 | 3.3 | 7.7 | |
| AMD EPYC 7R13 | Llama-3-8B Attention Layer | 1,400 | 2.1 | 200 | 12.1 | 7.0 | 40.3 |
| Deepseek-R1 MoE Layer | 2,290 | 1.7 | 330 | 2.3 | 6.9 | 9.4 | |
| FLUX Attention Layer | 2,460 | 1.5 | 230 | 3.1 | 10.7 | 22.1 | |
| FLUX Convolution Layer | 2,520 | 1.3 | 470 | 4.8 | 5.4 | 19.6 | |
| Llama-4-Scout MLP Layer | 510 | 6.4 | 100 | 10.2 | 5.1 | 8.1 | |
| Apple M2 Pro | Llama-3-8B Attention Layer | 1,010 | 3.3 | 190 | 9.7 | 5.3 | 15.6 |
| Deepseek-R1 MoE Layer | 1,040 | 2.8 | 230 | 4.8 | 4.5 | 7.8 | |
| FLUX Attention Layer | 270 | 2.1 | 50 | 3.7 | 5.4 | 9.5 | |
| FLUX Convolution Layer | 2,260 | 1.5 | 510 | 5.5 | 4.4 | 16.2 | |
| Llama-4-Scout MLP Layer | 2,460 | 2.2 | 440 | 3.4 | 5.6 | 8.6 | |
| Intel Core i9 | Llama-3-8B Attention Layer | 920 | 10.5 | 130 | 11.0 | 7.1 | 7.4 |
| Deepseek-R1 MoE Layer | 1,632 | 9.1 | 192 | 9.1 | 8.5 | 8.5 | |
| FLUX Attention Layer | 1,000 | 5.1 | 150 | 5.4 | 6.7 | 7.0 | |
| FLUX Convolution Layer | 400 | 2.3 | 72 | 2.3 | 5.6 | 5.6 | |
| Llama-4-Scout MLP Layer | 230 | 5.6 | 20 | 12.7 | 11.5 | 26.1 | |
| Intel Xeon E3 | Llama-3-8B Attention Layer | 2,760 | 3.9 | 320 | 5.8 | 8.6 | 12.8 |
| Deepseek-R1 MoE Layer | 1,000 | 3.7 | 180 | 4.4 | 5.6 | 6.6 | |
| FLUX Attention Layer | 1,340 | 1.4 | 450 | 3.4 | 3.0 | 7.1 | |
| FLUX Convolution Layer | 220 | 1.9 | 40 | 2.2 | 5.5 | 6.4 | |
| Llama-4-Scout MLP Layer | 1,200 | 2.0 | 300 | 6.1 | 4.0 | 12.2 | |
| Geomean | — | — | 2.7 | — | 5.0 | 5.8 | 10.8 |
@@
Speedup relative to baselines. LLM-Guided MCTS not only produces better code, but does so more aggressively and earlier in the search process. For example, on the DeepSeek-R1 MoE Layer, LLM-Guided MCTS achieves a 3.3 speedup over Evolutionary Search at 36 samples; on the Llama-4-Scout MLP Layer, LLM-Guided MCTS achieves a 9.3 speedup over Evolutionary Search at 20 samples. This trend, which shows strong initial gains followed by convergence, demonstrates that LLM-Guided MCTS quickly identifies high-performing regions of the search space, while Evolutionary Search’s uninformed search requires substantial exploration to reach similar quality.
@@
Operator-specific trends. We observe that certain operator types, such as matrix multiplication operations extracted from attention layers and MLP layer, exhibit sharper performance improvements. This is likely due to recurring structural patterns such as loop fusion, tiling, and vectorization, which pretrained LLMs can more readily recognize and exploit. Convolutional operators, by contrast, expose a broader and less regular transformation space. Nonetheless, Reasoning Compiler consistently matches or exceeds baseline performance with fewer samples, underscoring its effectiveness across diverse operator characteristics.
Sample efficiency across hardware platforms. As shown in Table 1, Reasoning Compiler demonstrates superior sample efficiency compared to TVM across five hardware platforms on five benchmarks. We define sample efficiency as the speedup achieved per sample (). On average, across all 25 platform-operator pairs, Reasoning Compiler achieves a 5.0 speedup using 5.8 fewer samples, resulting in a 10.8 improvement in sample efficiency. The performance gains are particularly significant for compute-intensive workloads. For instance, for the Llama-3-8B Attention Layer benchmark on AMD EPYC 7R13, Reasoning Compiler achieved a 12.1 speedup in just 200 samples, while TVM required 1,400 samples to reach a 2.1 speedup. This represents a 7.0 sample reduction and a 40.3 sample efficiency gain. On Intel Core i9, Reasoning Compiler often matches or exceeds TVM’s peak with fewer trials: on the Llama-4-Scout MLP Layer benchmark, Reasoning Compiler used 11.5 fewer samples for a 26.1 efficiency gain.
| Hardware Platform | TVM | Reasoning Compiler | Improvement | |||
| # Samples | Speedup | # Samples | Speedup | Sample Reduction | Sample Efficiency Gain | |
| Amazon Graviton2 | 4,560 | 3.7 | 1,440 | 5.1 | 3.2 | 4.4 |
| AMD EPYC 7R13 | 410 | 2.0 | 140 | 2.2 | 2.9 | 3.2 |
| Apple M2 Pro | 4,820 | 2.2 | 1,770 | 3.9 | 2.7 | 4.8 |
| Intel Core i9 | 3,800 | 2.2 | 720 | 4.9 | 5.3 | 11.8 |
| Intel Xeon E3 | 4,640 | 5.0 | 670 | 5.0 | 6.9 | 6.9 |
| Geomean | — | 2.8 | — | 4.0 | 3.9 | 5.6 |
End-to-end sample efficiency. For the end-to-end Llama-3-8B benchmark across the five hardware platforms in Table 2, Reasoning Compiler’s sample efficiency improvement over TVM ranges from 3.2 on AMD EPYC to 11.8 on Intel Core i9. End-to-end speedups range from 2.2 on AMD EPYC to 5.1 on Amazon Graviton2. Reasoning Compiler consistently achieves significantly higher speedups: using 3.9 fewer samples, it achieves a 4.0 speedup and yields a 5.6 geometric-mean sample efficiency improvement over TVM.
@@
Implications. These findings reinforce our core thesis: compiler optimization should be cast as a structured decision process, enriched by prior knowledge and contextual reasoning. Our integration of LLMs into Monte Carlo tree search results in a strategically guided and sample-efficient search, particularly valuable in scenarios with constrained tuning budgets. By generating performant code with orders-of-magnitude fewer samples, our framework offers both practical deployment advantages and a compelling alternative to conventional, sample-inefficient compilation pipelines.
@@ 4.3.1 Impact of LLM Choice and Reasoning Strategy
To better understand the contributions of different components in our approach, we conduct an ablation study focused on the effects of LLM selection and reasoning modality. Figure 4(a) shows the relative speedup over unoptimized code as a function of the number of schedule samples evaluated by LLM-Guided MCTS on the Llama-3-8B Attention Layer benchmark using a range of LLM models for API calls. The x-axis indicates the cumulative number of schedules explored, while the y-axis shows the best speedup achieved so far. This setup enables us to directly compare how effectively various LLMs leverage contextual information to guide the search. The general trend of the results supports our central claim: compiler optimization benefits from goal-directed, context-aware reasoning in terms of sample efficiency. Below, we discuss the specific behaviors that exemplify different reasoning strategies.
Large instruction-tuned Llama3.3 (70B) achieves exceptional sample efficiency. The instruction-tuned Llama3.3-70B model rapidly attains near-optimal performance, reaching a 9.69 speedup after only 36 samples, roughly 86% of the GPT-4o mini’s maximum speedup but with less than 6% of its sampling budget. This corresponds to an approximately 15 improvement in sample efficiency. Instruction tuning also significantly improves the ability of LLMs to generate domain-specific, context-aware transformation proposals. The consistent performance advantage of instruction-tuned models over untuned counterparts of comparable size confirms that semantic task alignment, combined with sufficient model capacity, synergistically enhances the effectiveness of sequential context reasoning in guiding compiler optimizations.
@@
Lower-parameter models also achieve high sample efficiency. Despite their reduced scale, smaller models still produce notable speedups relative to the untuned baseline. For example, Llama3.1-Instruct (8B) reaches a 5.87 speedup, and DeepSeek-R1-Distill-Qwen (7B) achieves a 4.86 speedup at just 36 samples. When compared to the widely used Evolutionary Search strategy, which requires around 72 samples to achieve a 7.0 speedup and fails to reach comparable performance for tuning the DeepSeek-R1 MoE Layer even after 3000 samples, these smaller models consistently outperform. LLM-Guided MCTS with lower-parameter models achieves at least twice the sample efficiency of Evolutionary Search, making them well-suited for efficient compiler optimization in local or edge deployments.
Open-source models match proprietary models in performance. Our results demonstrate that open-source LLMs, when adequately scaled and instruction-tuned, match or exceed the performance of proprietary baselines such as GPT-4o mini. This underscores the broad applicability of our approach and its independence from proprietary data or architectures, enabling widespread adoption of context-aware, LLM-guided compiler optimization.
@@ 4.3.2 Impact of Historical Trace Depth on Optimization Efficiency
Figure 4(b) presents the relative speedup over unoptimized code as a function of the number of schedule samples evaluated by LLM-Guided MCTS on the Llama-3-8B Attention Layer benchmark. Using a deeper historical trace (see Figure 1) in the prompt (parent + grandparent + great-grandparent) leads to faster convergence compared to the shallower trace (parent + grandparent). For example, at 36 samples, the deeper trace achieves a speedup of approximately 7.13, slightly surpassing the 7.08 of the shallower trace. However, by 72 samples, the deeper trace saturates at 11.36 speedup, while the shallower trace reaches only 8.38, requiring many more samples (around 579) to approach 11.3 performance. This demonstrates that including longer historical context enables the LLM to better capture dependencies and synergies in transformation sequences, resulting in more sample-efficient and goal-directed exploration, validating the advantage of context-aware reasoning.
@@ 5 Related Work
Superoptimization. While our high-level goal of discovering highly efficient program variants shares motivation with the superoptimization literature, our formulation and tractability differ substantially. Superoptimization aims to find the globally optimal instruction-level program, typically via enumerative [superoptimization:asplos:1987, peephole-superoptimizer:asplos:2006], symbolic [denali:pldi:2002], or stochastic [stoke:asplos:2013] search over low-level assembly variants; hybrid [lens:asplos:2016] and neural [learning-superoptimization:iclr:2017] approaches have also been explored. STOKE [stoke:asplos:2013] showed that high-quality programs often reside in low-probability regions and made the leap to use randomized search (MCMC). Neural compilers followed suit and relied on evolutionary search or simulated annealing algorithms [tensor-comprehensions:arxiv:2018, autotvm:neurips:2020, ansor:osdi:2020, metaschedule:neurips:2022]. In contrast, Reasoning Compiler treats optimization as a planning problem that leverages MCTS to reason contextually about dependencies among transformations over structured intermediate representations.
ML-Based Autotuning. Autotuning frameworks optimize performance-critical parameters (e.g., loop tile sizes, phase orderings, memory layouts) using a variety of ML-based techniques, including linear models [pred-unroll-factor:cgo:2005, micomp:taco:2017], tree-based methods [auto-construct-inlining-heuristic:cgo:2013, protuner:arxiv:2020], Bayesian networks [cobayn:taco:2016, baco:asplos:2023], evolutionary algorithms [auto-construct-inlining-heuristic:cgo:2013, mlgo:arxiv:2021, kmeans-auto-tuning:cgo:2025], clustering [micomp:taco:2017, kmeans-auto-tuning:cgo:2025], and reinforcement learning [chameleon:iclr:2020, autophase:mlsys:2020, mlgo:arxiv:2021, kmeans-auto-tuning:cgo:2025]. Reasoning Compiler shares the same goal of performance-driven parameter selection, but distinguishes itself by combining LLM-based contextual reasoning with structured search (via MCTS) to explore transformation sequences in a history- and structure-aware manner.
Machine Learning for Neural Compilation. Machine learning has been extensively applied to optimize neural network inference pipelines, including high-level graph-level optimizations [dl-with-dynamic-graph:arxiv:2017, taso:sosp:2019, neocpu:atc:2019, go:neurips:2020, ios:mlsys:2021, fusionstitching:arxiv:2021, equality-saturation-tensor-graph-superopt:mlsys:2021, apollo:mlsys:2022] and low-level code generation [tensor-comprehensions:arxiv:2018, tvm, tiramisu:cgo:2019, fireiron:pact:2020, unit:cgo:2021, dse-cnn:asplos:2021, deepcuts:pldi:2021, hidet:asplos:2023]. Systems such as TVM/Ansor [tvm, autotvm:neurips:2020, ansor:osdi:2020] and FlexTensor [flextensor:asplos:2020] employ learned cost models and evolutionary strategies to navigate large configuration spaces. While these approaches are highly effective at tuning tensor programs, they typically focus on local parameter optimization or rely on domain-specific heuristics. Reasoning Compiler moves beyond these works by introducing contextual reasoning through LLMs, enabling the system to reason over transformation history, structural changes, and performance trends, an approach not explored in prior neural compilation work.
LLMs for Code Reasoning and Optimization. LLMs have demonstrated capabilities in code generation [codellama, starcoder, magicoder, codegeex, codet5, deepseekcoder], fuzzing [llm-fuzzing], bug repair [prog-repair-llm], and even high-level optimization [pie:iclr:2024]. Recent work has explored the use of LLMs to generate phase orderings or perform disassembly [llm-compiler-optimization:arxiv:2023, llm-compiler:cc:2025]. Reasoning Compiler advances these approaches by embedding an LLM in a structured decision loop, leveraging it for context-aware reasoning within a grounded search process.